A firm has the following revenue and cost functions.



TR = 120 Q – Q2



TC = 1/2 Q2 +30 Q + 10



Determine the quantity level at which the firm maximizes its total profit. (Hint: use marginal revenue = marginal cost rule)

1
Expert's answer
2022-02-08T14:09:39-0500

Solution:

Derive Marginal Revenue = Marginal Cost

MR = "\frac{\partial TR} {\partial Q} = 120 - 2Q"

MC = "\frac{\partial TC} {\partial Q} = Q + 30"

Set MR = MC

120 – 2Q = Q + 30

120 – 30 = Q + 2Q

90 = 3Q

Q = 30

The quantity level at which the firm maximizes its total profit = 30 units